OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3

OpenStudy (anonymous):

Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sid

OpenStudy (anonymous):

This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (anonymous):

SOMEONE HELP MEEEEEE

OpenStudy (ghazi):

rearrange the equation..... 3x-2x=2-4-1= -3...proved....

OpenStudy (anonymous):

so what would my answer be ?

OpenStudy (ghazi):

it would be X=-3

OpenStudy (anonymous):

how would i write that out ?

OpenStudy (ghazi):

well can you solve this 3x-2x= ?

OpenStudy (ghazi):

\[3x-2x= 2-4-1\] next step would be \[x=-3 \] that's it...you're done

OpenStudy (ghazi):

hope this helps

OpenStudy (anonymous):

thank you ! :)

OpenStudy (ghazi):

:)